OpenStudy (anonymous):

Find the derivative of f(x)=3/sqrt(x^3-2)

OpenStudy (anonymous):

\[\Large f(x)={3 \over \sqrt{x^3-2}}\] correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Alright do you know hot to do the chain rule?

OpenStudy (triciaal):

@jamie1522 do you know how?

OpenStudy (anonymous):

Yea sort of

OpenStudy (triciaal):

dy/dx= vdu/dx-udv/dx all over v2

OpenStudy (anonymous):

Ok, well it just means that you have to take the derivative of everything in your "outer function" and multiply it by the derivative of your "inner function" \[\Large {dy \over dx}f(g(x))=f \prime (g(x)) \cdot g \prime(x)\]

OpenStudy (anonymous):

Now f(x)= \[\Large {3 \over \sqrt{x}}\] and g(x)= \[\Large x^3-2\] Now I want you to first off, take the derivative of f(x) and g(x) and then we'll go from there

OpenStudy (triciaal):

with my approach u = numerator and v = denominator for this problem u = 3 and v = (x^3-2)^1/2

OpenStudy (triciaal):

continue with @doulikepiecauseidont since he/she responded first. the answer should be the same

OpenStudy (anonymous):

\[f'(x)=-\frac{ 3 }{ 2x ^{\frac{ 3 }{ 2 }} }\] and \[g'(x)=3x ^{2}\]

OpenStudy (anonymous):

Alright now with the equation for f'(x), plug in g(x) which is x^3-2 then you want to multiply that by g'(x) which is 3x^2

OpenStudy (anonymous):

So that is \[\frac{ -9x ^{2} }{ 2(3x ^{\frac{ 7 }{ 2 }}) }\] correct?

OpenStudy (anonymous):

Not exactly, you want to plug in 3x^2-2 into f'(x) wherever there is an x so It would be \[\Large f'(g(x))=-\frac{3}{2(3x^2-2)^\frac{3}{2}}\]

OpenStudy (anonymous):

Now this will be mulitplied by our g'(x) which is 3x^2

OpenStudy (anonymous):

Wait why is it 3x^2-2? I thought you said plug in g(x)? which is x^3-2

OpenStudy (anonymous):

Sorry, got my exponents mixed up, you are correct it would be x^3-2

OpenStudy (anonymous):

Alright. So would it be \[\frac{ -9x ^{2} }{ 2(x ^{\frac{ 9 }{ 2 }}2^{\frac{ 5 }{ 2 }} )} \] ?

OpenStudy (anonymous):

\[\Large f'(g(x))=-\frac{3}{2(3x^2-2)^\frac{3}{2}} \] would be multiplied by g'(x)

OpenStudy (anonymous):

3x^2

OpenStudy (anonymous):

\[\Large f'(g(x))=-\frac{9x^2}{2(3x^2-2)^\frac{3}{2}}\]

OpenStudy (anonymous):

Is that the final answer?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

\[\Large {dy \over dx}f(g(x))=f \prime (g(x)) \cdot g \prime(x)\]

OpenStudy (anonymous):

You have to be able to identify the outer function (in our case it was the sqrt function) And the inner function (x^3-2)

OpenStudy (anonymous):

outer=f(x) inner=g(x)

OpenStudy (anonymous):

Okay thank you so much for your help!

OpenStudy (anonymous):

Yw

OpenStudy (triciaal):

@jamie1522 was that the correct answer? @doulikepiecauseidont can you check my answer was -9 x^2 / 2(x^3-2)^1/2

OpenStudy (anonymous):

Your exponent is wrong, remember that (x^3-2) was already raised to the 1/2 power (that's where the sqrt came from) and then when you apply the power rule (since it's in the denominator and is then like being raised to the -1/2 power) you get the - on the outside and the 2 in the denominatior times the function raised to the -3/2 or 1/(x)^3/2

OpenStudy (triciaal):

my exponent cannot be wrong 1/2 - 1 = -1/2 when you differentiate  maybe yours can be simplified or you made a mistake look again, that's not the only thing you have different